OpenStudy (anonymous):

If a phone company charges $0.35 for the first minute and $0.25 for each additional minute for a long-distance call, how long did you talk if the phone call cost $7.85? You talked for how many minutes? Enter your answer as a whole or decimal number.

OpenStudy (anonymous):

0,35+0,25(x-1)=7,85

OpenStudy (anonymous):

31 minutes

OpenStudy (johnweldon1993):

so .35m + .25a = 7.85 .35(1) + .25a = 7.85 .35 + .25a = 7.85 .25a = 7.50 a = 30 minutes of additional plus the original 1 of normal talking...so 31 minutes

OpenStudy (anonymous):

\[\frac{7.85-0.35}{0.25}+1=31 \]